# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

### v2

@@ 5.1 GB space charge layers at equilibrium
To demonstrate the capability of the phase-field model at thermodynamically equilibrium state, one-dimensional phase-field simulations are conducted and compared with the sharp interface predictions [30]. The semi-finite simulation domain has the size of 780 nm, with the core located at the left boundary. The grain boundary thickness is prescribed to be 0.78 nm [30]. The domain is discretized into a mesh consisting of 2000 elements. The simulation temperature is set at 600 K, and the relative permittivity $\epsilon_{\text{r}}$ of strontium titanate (STO) is specified as 160, according to [30], which is calculated via $\epsilon_{\text{r}}=9\times 10^{4}/(T-35)$ as mentioned in section 3.1. The interface energy, molar volume and interface mobility of STO are 0.6 $\text{J}\text{m}^{-2}$, 3.6×$\times 10^{-5}$ $\text{m}^{3}\text{mol}^{-1}$ and 1×$\times 10^{-7}$ $\text{m}^{3}\text{J}^{-1}\text{s}^{-1}$, respectively [38]. The relative tolerance for convergence is set to $1\times 10^{-8}$.
5.1.1 Benchmark case 1: The MS Model
By fixing a uniform dopant concentration, the proposed phase-field model can be used to simulate the MS case. The results are given in Fig. 2. The initial concentrations of oxygen vacancies and dopant far from the core (denoted as $c_{\text{dop},\text{b}}(\infty)$ and $c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$, respectively) are $7\times 10^{24}\text{ m}^{-3}$ and $3.5\times 10^{24}\text{ m}^{-3}$ [30]. The number of available sites of oxygen vacancy in bulk phase and GB core are $N_{\text{V}_{\text{O}}^{..},\text{b}}=5.12\times 10^{28}$ $\text{m}^{-3}$ and $N_{\text{V}_{\text{O}}^{..},\text{c}}=3\times 10^{26}$ $\text{m}^{-3}$, respectively[30, 49]. The grain boundary potential $\Phi_{0}$ is plotted as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$ in Fig. 2(a). Due to the right boundary being grounded, the grain boundary potential is directly obtained at the interface position when $\eta=0.5$, as defined in the phase-field method. For $\Delta g_{\text{V}_{\text{O}}^{..}}$ larger than -1.25 eV, the phase-field results exhibit good agreement with analytical calculations. As $\Delta g_{\text{V}_{\text{O}}^{..}}$ becomes more negative, particularly at -1.5 eV and -2.0 eV, the available oxygen vacancy sites are completely exhausted. Therefore, the concentrations of oxygen vacancies in the core are close to $N_{\text{V}_{\text{O}}^{..},\text{c}}$. Although the phase-field results demonstrate convergent behavior, a discrepancy between the phase-field results and the analytical solution emerges. To explain the aforementioned discrepancy, we plot the concentration of oxygen vacancies in the core as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$ in Fig. 2(b). Similar to the grain boundary potential, $c_{\text{V}_{\text{O}}^{..}}$ is also extracted at $\eta=0.5$. We observe that the phase-field results coincide with the analytical predictions, even when $\Delta g_{\text{V}_{\text{O}}^{..}}=-2.0$ eV. Next, we examine the concentration of oxygen vacancy in the bulk in terms of $x$. Here, we plot the phase-field results for $\Delta g_{\text{V}_{\text{O}}^{..}}$=-0.5 eV, -1.0 eV and -1.5 eV in Fig. 2 (c), (d) and (e) in the range of 0 to 50 nm, which includes half of the interface region. For the analytical solution, we shift it to the right by a distance of $w_{\text{c}}/2$ to match the bulk phase in the phase-field results. When ${\Delta g_{\text{V}_{\text{O}}^{..}}}=-0.5$ eV, $c_{\text{V}_{\text{O}}^{..},\text{b}}$ obtained by phase-field simulations agrees well with analytical predictions as shown in Fig. 2(c1). However, in Fig. 2(d1) and (e1), for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ and $-2.0$ eV, the phase-field results produce higher $c_{\text{V}_{\text{O}}^{..},\text{b}}$. This comparison indicates that less oxygen vacancy is segregated in the core in the phase-field simulations when choosing more negative $\Delta g_{\text{V}_{\text{O}}^{..}}$. Additionally, we plot the area density of the net charge, $Q_{\mathrm{net}}=\int_{-w_{\text{c}}/2}^{w_{\text{c}}/2}(z{{}_{\text{dop}}}c_{{}_{\text{dop},\text{c}}}+z_{{}_{\text{V}_{\text{O}}^{..}}}c_{{}_{\text{V}_{\text{O}}^{..},\text{c}}})\text{d}x$, in the core in Fig. 2(f). Less net charge is reproduced via phase-field simulation when $\Delta g_{\text{V}_{\text{O}}^{..}}$ is smaller. In short, this discrepancy is attributed to the diffuse interface in the phase-field simulations. Due to its finite interface thickness and smooth interface characteristics, the core is less positively charged compared to analytical solutions. Consequently, the electrostatic potential profile obtained via phase-field simulations decays faster than analytical calculations as shown in Fig. 2(c2), (d2) and (e2). Up to this point, we clearly observe a good agreement between the oxygen vacancy concentrations in the GB core and bulk phase as predicted by the phase-field results and the sharp interface predictions. The different approaches to handling interfaces in the phase-field model and the sharp interface method can result in discrepancies when the available densities of point defects are fully occupied.
@@ 5.1.1 Benchmark case 1: The MS Model
Figure 2: Comparison of phase-field (PF) simulation outcomes with those of the sharp interface (SI) model predictions obtained via FEM and taken from Ref. [30] in the MS model. Figure (a) illustrates the grain boundary potential as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$. Figure (b) depicts the relationship between the oxygen vacancy concentration and $\Delta g_{\text{V}_{\text{O}}^{..}}$. Figures (c), (d), and (e) present the concentration distribution of dopant and the electrostatic potential distribution for different oxygen vacancy segregation energy. Figure (f) demonstrates the net area density of charge in the core as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$.
5.1.2 Benchmark case 2: The GC Model
By setting the electrochemical potentials of the acceptor dopant and of the oxygen vacancies to be constant, the presented phase-field model is then reduced to the GC model, in order to reproduce the equilibrium SCL formation in the GC case. The number of available sites of dopant and oxygen vacancy in bulk and GB core are $N_{\text{V}_{\text{O}}^{..},\text{c}}=1\times 10^{27}$ $\text{m}^{-3}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}=5.12\times 10^{28}$ $\text{m}^{-3}$, $N_{\text{dop},\text{c}}=1.68\times 10^{27}$ $\text{m}^{-3}$ and $N_{\text{dop},\text{b}}=1.69\times 10^{28}$ $\text{m}^{-3}$, respectively [30, 49]. The Debye lengths $l_{\text{D}}=\sqrt{\frac{\epsilon_{0}\epsilon_{\text{r}}k_{\text{B}}T}{2e^{2}c_{\text{dop},\text{b}}(\infty)}}$ in both MS and GC models are 5.71 nm. In Fig. 3(a), the grain boundary potentials obtained via the phase-field simulations are compared with the sharp interface predictions. $\Delta g_{\text{V}_{\text{O}}^{..}}$ is fixed at -1.0 eV here.
When $\Delta g_{\text{dop}}$ is positive, dopant cations are rejected from the core, thereby causing a more pronounced grain boundary potential. Conversely, as $\Delta g_{\text{dop}}$ becomes negative, dopant segregation to the core occurs, reducing the net charge and consequently decreasing the grain boundary potential. Moreover, as $\Delta g_{\text{dop}}$ becomes sufficiently negative, the available number sites for dopant segregating within the core become scarce. Therefore, the grain boundary potential profile is flat when $\Delta g_{\text{dop}}$ is positive or sufficiently negative. The comparison indicates the phase-field results agree with the predictions of analytical solution. The discrepancy is smaller than 7%. The concentrations of vacancies and acceptor dopants in the core as a function of $\Delta g_{\text{dop}}$ are plotted in Fig. 3(b).
In Fig. 3(c), (d), and (e), we present the concentration profiles of oxygen vacancies and dopant, as well as the distribution of electrostatic potential in the bulk phase for $\Delta g_{\text{dop}}$ = 0 eV, -0.125 eV and -0.375 eV, respectively. Due to symmetry, we only depict half of the SCL in one bulk phase.
At $\Delta g_{\text{dop}}$ = 0 eV, phase-field results predict the formation of a dopant accumulation zone near the core. The dopant concentration in the core can be much lower than in the accumulation zone, leading to the formation of a nonmonotonic curve as shown in Fig. 3(c1). Contrarily, in Virkant’s findings [38], the expected accumulation zone at $\Delta g_{\text{dop}}$=0 eV is replaced by a depletion zone. As $\Delta g_{\text{dop}}$ decreases to -0.125 eV, dopant concentration in the core increases significantly, causing the formation of monotonic concentration profile. Further decreasing $\Delta g_{\text{dop}}$ to -0.375 eV, the dopant concentration in the core increases to $N_{\text{dop},\text{b}}$. It is evident that the accumulation zone becomes lower than 0 eV, more exactly -0.125 eV. Additionally, near the core, the concentration profile of dopant exhibits a distinct kink, indicating a notable change in the gradient.
The oxygen vacancy concentration profiles of are plotted in Fig. 3(c2), (d2) and (e2). Depletion appears for all $\Delta g_{\text{dop}}$. Notably, as $\Delta g_{\text{dop}}$ decreases, there is an increase in the concentrations within these depletion zones. In Fig. 3(c3), (d3) and (e3), we present a comparison of the electrostatic potential across the SCL with analytical solutions. The phase-field simulation results agree well with these analytical predictions. Decreasing $\Delta g_{\text{dop}}$ causes lower electrostatic potential distribution near the core. Moreover, we also present the net charge area density $Q_{\mathrm{net}}$ in Fig. 3(f). The trend of the curve obtained via phase-field results agrees with the analytical predictions. Analogous to the MS case, phase-field results are smaller than analytical predictions.

### v3

@@ 5.1 GB space charge layers at equilibrium
To examine the influence of various defect parameters on SCL formation, 1D phase-field simulations are conducted for a bicrystal configuration at temperatures of $T=600$ K and $T=1623$ K. At 600 K, the distribution of acceptor dopant is frozen, while that of the oxygen vacancy achieves chemical equilibrium (MS model). At 1623 K, the diffusion of acceptor dopant becomes active, enabling both the acceptor dopant and oxygen vacancy to achieve electrochemical equilibrium, referred to as the GC model. Furthermore, we chose the physical GB width (0.78 nm de2009formation ) in the phase-field simulations. This choice can decrease the abnormal interface effects raised by an artificially enlarged interface thickness wang2020modeling , wang2021quantitative . The defect chemistry parameters used for the phase-field simulations of SCLs are summarized in Table 3.
We summarize the numerical settings employed in the FEM implementation. The simulation domain is one-dimensional and symmetric, spanning from x=390x=-390 nm to $x=390$ nm, with the GB located at $x=0$ nm. The domain is discretized into 10,000 elements with a uniform element size of 0.078 nm. For the electrostatic potential, a Neumann boundary condition is applied at the right boundary, while the left boundary is grounded. Neumann boundary conditions are also imposed on the concentration field at both boundaries. Each simulation is executed on a single CPU node with 16 cores and 8 gigabyte of memory, requiring approximately 5 hours of computation time.
Table 3: The defect chemistry parameters for the 1D phase-field simulations in MS and GC models to reproduce equilibrium SCLs. The definition of these parameters are described in Table 1.
Parameters Value in MS model de2009formation Value in GC model de2009formation
TT/K 600 1623
aa/nm 0.39 0.39
$\omega_{c}$/nm 0.78 0.78
$\epsilon_{r}$ 160 56
$c_{\text{dop},\text{b}}(\infty)$/ $\text{m}^{-3}$ $7\times 10^{25}$ $7\times 10^{25}$
$c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$/ $\text{m}^{-3}$ $3.5\times 10^{25}$ $3.5\times 10^{25}$
$l_{\text{D}}$/nm 1.81 1.76
$\Delta g_{\text{V}_{\text{O}}^{..}}$/eV -1.5 $\sim$ 0 -1.5
$\Delta g_{\text{dop}}$/eV - -1.0 $\sim$ 0.5
$N_{\text{V}_{\text{O}}^{..},\text{b}}$/ $\text{m}^{-3}$ $5.1\times 10^{28}$ $5.1\times 10^{28}$
$N_{\text{dop},\text{b}}$/ $\text{m}^{-3}$ - $1.68\times 10^{28}$
$N_{\text{V}_{\text{O}}^{..},\text{c}}$/ $\text{m}^{-3}$ $5\times 10^{26}$ $\sim$ $1\times 10^{27}$ $5\times 10^{26}$ $\sim$ $1\times 10^{27}$
$N_{\text{dop},\text{c}}$/ $\text{m}^{-3}$ - $1.68\times 10^{26}$ $\sim$ $1.68\times 10^{27}$ de2009formation
5.1.1 Benchmark case 1: MS Model
By prescribing a uniform dopant concentration, the proposed phase-field model is employed to simulate the MS case. Phase-field simulations of SCLs with varying available number densities in the MS model are systematically benchmarked against the sharp interface (SI) calculations of GB||SCL configurations, as described in Section 2. The results of this comparison are presented in Fig. 2. The phase-field results are represented by solid lines, while the SI calculations are denoted by cross symbols. The orange, blue, and purple colors correspond to different available site densities: orange for $N_{\text{V}_{\text{O}}^{..},\text{c}}=1\times 10^{27}\text{m}^{-3}$, blue for $N_{\text{V}_{\text{O}}^{..},\text{c}}=7.5\times 10^{26}\text{m}^{-3}$, and purple for $N_{\text{V}_{\text{O}}^{..},\text{c}}=5\times 10^{26}\text{m}^{-3}$.
Fig. 2(a) and (b) illustrate the GB potentials ($\Phi_{0}$) and oxygen vacancy concentrations at the GB core ($c_{\text{V}_{\text{O}}^{..},\text{c}}$) for various segregation energies and available oxygen vacancy number densities. It should be clarified that $\Phi_{0}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$ are directly obtained at the interface position, defined by $\eta=0.5$ in the phase-field method. An increase in $|\Delta g_{\text{V}_{\text{O}}^{..}}|$ ($\Delta g_{\text{V}_{\text{O}}^{..}}<0$) promotes greater oxygen vacancy segregation at the GB core, thereby enhancing the GB potential. However, the GB potential does not increase indefinitely. At sufficiently large $|\Delta g_{\text{V}_{\text{O}}^{..}}|$, when all available sites for oxygen vacancies are fully occupied, the GB potential reaches a plateau, as does the oxygen vacancy concentration in the GB core ($c_{\text{V}_{\text{O}}^{..},\text{c}}\approx N_{\text{V}_{\text{O}}^{..},\text{c}}$, see Fig. 2(b)). The available number density of oxygen vacancy sites ($N_{\text{V}_{\text{O}}^{..},\text{c}}$) plays a critical role in determining the plateau value. For instance, when $N_{\text{V}_{\text{O}}^{..},\text{c}}$ increases from $5\times 10^{26}\text{ m}^{-3}$ to $7.5\times 10^{26}\text{ m}^{-3}$, the plateau GB potential rises from 0.12 V to 0.24 V in the phase-field results and from 0.13 V to 0.27 V in the sharp interface (SI) calculations. However, when $N_{\text{V}_{\text{O}}^{..},\text{c}}$ is as high as $1\times 10^{27}\text{ m}^{-3}$, this plateau behavior becomes less pronounced, as no clear convergence of the GB potential is observed in both the SI calculations and phase-field results. Although phase-field results exhibit the same trend as predicted by SI calculations, discrepancies in GB potentials are still observed in the large $|\Delta g_{\text{V}_{\text{O}}^{..}}|$ region.
To further investigate the GB core, we plot the area density of charge in the GB core, $Q_{\text{c}}=\int_{-w_{c}/2}^{w_{c}/2}e(z_{\text{dop}}c_{\text{dop},\text{c}}+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{c}})\text{d}x$, obtained from phase-field results and SI calculations, as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$ for different $N_{\text{V}_{\text{O}}^{..},\text{c}}$ in Fig. 2(c). Here, $Q_{\text{c}}$ represents the boundary condition in the SI calculations, as defined in Eq. 17, which plays a critical role in determining the GB potential and defect concentration in GB core. When $|\Delta g_{\text{V}_{\text{O}}^{..}}|$ is low, the phase-field results align well with the SI calculations. However, discrepancies emerge in the high $|\Delta g_{\text{V}_{\text{O}}^{..}}|$ region, where phase-field simulations reproduce lower area charge densities. These discrepancies are likely due to the diffuse interface in phase-field simulations. In the SI model, the defect concentrations between the GB core and the bulk phase are discontinuous, with defect concentrations assumed to be constant across the GB core. Conversely, the smooth interface characteristics inherent to phase-field simulations result in a less positively charged GB core compared to the SI calculations when the same GB width, $w_{c}=0.78$ nm, is employed.
In addition to the GB core, we also compare the distributions of electrostatic potentials and oxygen vacancy concentrations in the bulk phase for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ values, as shown in Fig. 2(d) to (f). Due to the symmetric bicrystal configuration, only half of the interface region is included in the analysis. In Fig. 2 (d1), (e1), and (f1), depletion zones of oxygen vacancy are observed near the GB core. Larger $N_{\text{V}_{\text{O}}^{..},\text{c}}$ values result in more pronounced depletion of oxygen vacancy. Overall, the phase-field results show good agreement with the SI calculations. n Fig. 2(d1), (e1), and (f1), depletion zones of oxygen vacancy are observed near the GB core. Larger $N_{\text{V}_{\text{O}}^{..},\text{c}}$ values lead to more pronounced depletion of oxygen vacancy. Overall, the phase-field results align well with the SI calculations. However, discrepancies are observed near the GB core, where oxygen vacancy concentrations decrease in the SI calculations but increase in the phase-field results (see Fig. 2(e1) and (e2) in the range of [0, 2 nm]). These differences stem from the contrasting assumptions used to connect the GB core and bulk phase in the phase-field model and the discontinuous SI model, as discussed earlier. In Fig. 2 (d2), (e2), and (f2), larger $N_{\text{V}_{\text{O}}^{..},\text{c}}$ valules lead to higher electrostatic potential and longer electrostatic decay lengths. The phase-field simulations is capable of capturing these trends and nicely reproducing the SI predictions.
@@ 5.1.1 Benchmark case 1: MS Model
Figure 2: (a) Comparison of phase-field (PF) simulation outcomes with those of the abrupt GB||SCL sharp interface (SI) predictions with different available number density of oxygen vacancy, such as $N_{\text{V}_{\text{O}}^{..},\text{b}}=5\times 10^{26}\text{m}^{-3}$, $7.5\times 10^{26}\text{m}^{-3}$ and $1\times 10^{27}\text{m}^{-3}$. Figure (a), (b) and (c) illustrates the GB potential, GB concentration of oxygen vacancy and area density of charge as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$. Figure (d), (e) and (f) depicts the spatial distribution of oxygen vacancy concentration and electrostatic potential in bulk phase.
5.1.2 Benchmark case 2: GC Model
By fixing the electrochemical potentials of the acceptor dopant and oxygen vacancies as constants, the proposed phase-field model reduces to the GC model, enabling the reproduction of equilibrium SCL formation in the GC case. In this framework, the effects of the segregation energy of the acceptor dopant ($\Delta g_{\text{dop}}$) and the available site densities of oxygen vacancies and acceptor dopants in the GB core ($N_{\text{dop},\text{c}}$ and $N_{\text{V}_{\text{O}}^{..},\text{c}}$) on SCL formation are investigated and compared with SI calculations, as illustrated in Fig. 3. In both phase-field and SI results, the segregation energy of oxygen vacancy $\Delta g_{\text{V}_{\text{O}}^{..}}$ is set to 1.5 eV. For $N_{\text{V}_{\text{O}}^{..},\text{c}}=1\times 10^{27}\text{m}^{-3}$, phase-field results and SI predictions are shown using solid lines and cross symbols, respectively. For $N_{\text{V}_{\text{O}}^{..},\text{c}}=5\times 10^{26}\text{m}^{-3}$, dashed lines and triangle symbols are used. Additionally, different colors are employed to represent varying $N_{\text{dop},\text{c}}$: red for $N_{\text{dop},\text{c}}=1.68\times 10^{26}\text{m}^{-3}$, blue for $N_{\text{dop},\text{c}}=8.4\times 10^{26}\text{m}^{-3}$, and orange for $N_{\text{dop},\text{c}}=1.68\times 10^{27}\text{m}^{-3}$.
In Fig. 3(a), (b), and (c), the GB potential ($\Phi_{0}$), GB concentrations of oxygen vacancy ($c_{\text{V}_{\text{O}}^{..},\text{c}}$), and acceptor dopant ($c_{\text{dop},\text{c}}$) are presented as functions of $\Delta g_{\text{dop}}$. In Fig. 3(a), the GB potential curves exhibit two distinct plateaus. For positive $\Delta g_{\text{dop}}$, the GB potential reaches an upper plateau due to minimal dopant segregation in the GB core. In this case, the dopant concentration in the GB core can be 2 to 3 orders of magnitude lower than in the bulk phase (see Fig. 3(b) at positive $\Delta g_{\text{dop}}$ region), making it insufficient to compensate for the charge of the GB core induced by oxygen vacancy. As $\Delta g_{\text{dop}}$ becomes negative, the positively charged GB core is gradually compensated by the segregating dopant (see Fig. 3(b) at negative $\Delta g_{\text{dop}}$ region), leading to a reduction in GB potential. When all available sites for acceptor dopant are fully occupied, the GB potential reaches its lower plateau. Increasing $N_{\text{V}_{\text{O}}^{..},\text{c}}$ raises both the upper and lower plateaus simultaneously, with little impact on the shape of the GB potential curve. The difference between the upper and lower plateaus is primarily determined by $N_{\text{dop},\text{c}}$, with larger $N_{\text{dop},\text{c}}$ resulting in a greater plateau difference. For example, when $N_{\text{V}_{\text{O}}^{..},\text{c}}=5\times 10^{26}\text{m}^{-3}$, the plateau difference is 0.17 V for $N_{\text{dop},\text{c}}=1.68\times 10^{26}\text{m}^{-3}$ and increases to 1.5 V for $N_{\text{dop},\text{c}}=1.68\times 10^{27}\text{m}^{-3}$. In addition, very large $N_{\text{dop},\text{c}}$ might lead to negative GB potential as presented by the orange dashed line and triangle symbols when $\Delta g_{\text{dop}}$ is smaller than -0.8 eV in Fig. 3(a).
Adjusting $N_{\text{V}_{\text{O}}^{..},\text{c}}$ and $N_{\text{dop},\text{c}}$ has a pronounced effect on the GB concentrations of oxygen vacancy and acceptor dopant. When the available sites are not fully occupied, increasing $N_{\text{V}_{\text{O}}^{..},\text{c}}$ and $N_{\text{dop},\text{c}}$ results in greater dopant accumulation within the GB core for a given $\Delta g_{\text{dop}}$, as shown in Fig. 2(b). The segregation of dopant also influences the distribution of oxygen vacancy. In Fig. 2(c), a more negative $\Delta g_{\text{dop}}$ leads to a higher GB concentration of oxygen vacancy. Similar to the GB potential curves, the oxygen vacancy concentration curves exhibit two plateaus. An increase in $N_{\text{dop},\text{c}}$ amplifies the difference between these plateaus.
We also present the distributions of electrostatic potential, oxygen vacancy concentration, and acceptor dopant concentration in the bulk phase for different $N_{\text{dop},\text{c}}$ and $\Delta g_{\text{dop}}$ in Fig. 3(d), (e), and (f). Here, $N_{\text{V}_{\text{O}}^{..},\text{c}}$ is fixed at $5\times 10^{26}\text{m}^{-3}$. First, we observe a depletion zone for oxygen vacancy and an accumulation zone for acceptor dopant in the SCLs when the GB potential is positive, while the opposite behavior is observed for negative GB potential (see the orange dashed lines and triangle symbols in Fig. 3(f1), (f2), and (f3)). For $\Delta g_{\text{dop}}=0$ eV, increasing $N_{\text{dop},\text{c}}$ has minimal impact on the distributions of oxygen vacancy concentration, acceptor dopant concentration, and electrostatic potential in the bulk phase. As a result, the curves for $N_{\text{dop},\text{c}}=1.68\times 10^{26}\text{m}^{-3}$, $8.4\times 10^{26}\text{m}^{-3}$, and $1.68\times 10^{27}\text{m}^{-3}$ are nearly indistinguishable, as shown in Fig. 3(d1), (d2), and (d3). For $\Delta g_{\text{dop}}=-0.5$ eV and 1.0-1.0 eV, increasing $N_{\text{dop},\text{c}}$ raises the oxygen vacancy concentration in the depletion zone, reduces the acceptor dopant concentration in the accumulation zone, and decreases the electrostatic potential in SCLs. However, we still observe discrepancies between phase-field results and SI predictions near the GB core, which has been clarified in the MS case.
We compare the equilibrium SCL formation obtained from phase-field simulations with the results of Vikrant’s simulations for Fe-doped SrTiO3 (see Figure 1 in Ref. vikrant2020electrochemical ). In Vikrant’s study, a higher dopant concentration is used, i.e. 2% (in present work, dopant concentration is 0.42%), which is considered as non-dilute case and different grain boundary misorientation is considered. Their simulation temperature is set at 1623 K, and the acceptor dopant is considered mobile. They employed a negative segregation energy for oxygen vacancy (0.8-0.8 eV) and a zero segregation energy for dopant.
The major differences between the two studies are as follows: (i) In Vikrant’s results, the dopant concentration in the GB core is consistently higher than in the bulk phase for all GB misorientations. (ii) Their simulations also reveal a depletion zone for acceptor dopant near the GB core within the SCL. In the present work, a zero segregation energy for dopant results in significantly lower dopant concentrations in the GB core. As shown in Fig. 3(b), the dopant concentration in the GB core, $c_{\text{dop},\text{c}}$, is consistently smaller than the bulk concentration, $c_{\text{dop},\text{b}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$, for all cases where $\Delta g_{\text{dop}}=0$ eV. Furthermore, we observe an accumulation zone for acceptor dopant in the SCL region under positive electrostatic potentials, as depicted in Fig. 3(d2), (e2) and (f2).
We note that defect–defect interactions were included in Vikrant’s work and can influence space charge layer formation, as discussed in Refs. lund2021thermodynamically , mebane2015generalised. However, their model assumes that oxygen vacancies and dopants share the same sublattice, which affects the configurational entropy and consequently the electrochemical potential (see Section 3.3), particularly in the non-dilute regime where these quantities are sensitive to the underlying sublattice occupancy assumptions.
Thus far, we have successfully benchmarked the capabilities of the proposed phase-field model against the abrupt GB||SCL model in both the MS and GC cases. By comprehensively accounting for the segregation energies of various defects and partitioning the available site densities among different defects, diverse SCL formation behaviors have been accurately reproduced. As a result, the phase-field results show excellent agreement with SI predictions, not only within the GB core but also in the bulk phase. However, due to the assumptions of discontinuous GB and SCL in the SI model, two key discrepancies are observed. First, the GB potential in the GB core obtained from the phase-field model is smaller than that predicted by the SI model, particularly when the segregation energy is sufficiently large. Second, differences in defect concentrations are confined to the immediate vicinity of the GB core.